The perimeter of a rectangular garden is 27 feet. The length of the garden is twice the width. What is the length and width of the garden?
Answer by vicgonzerx(31)

perimeter formula for rectangular
P=2(l+w)
since the l=2w
P=27
so;
27=2(2w+w)
27=6w
w=4.5ft
L=2w
L=2(4.5)
L=9ft is the answer